QUESTION 5 5. Loan Interest Paid Dean is planning to purchase a new Nissan Altima which costs $26,865. He has saved up $3,200 to put toward it. He is thinking about getting a loan for it from his credit union, but he could also keep driving his beat-up old car for a while longer and save up cash for the new one. Possibility One: A Loan Dean qualifies for a 5-year loan from his credit union with a 6.4% APR. Assumw he uses the $3,200 as a down payment. You have already calculated the total he will pay for the car, including all the monthly payments and the down payment. Compare that to the sticker price of the car: how much of the total Dean spends on the car will be interest?
Answer #1

I = 6.4% / 12 = 0.5333% per month

t = 5 * 12 = 60 months

Loan amount = 26865 - 3200 = 23665

Monthly loan payment = 23665 * (A/P, 0.5333%,60)

= 23665 * 0.019519

= 461.92

Total interest payment = 461.92 * 60 - 23665 = 4050.25